OpenStudy (anonymous):

suppose the diameter of an electric cable is normally sistributed with mean 0.8 and standard deviation .02. a cable is considered defective if the diameter differs from the mean by more than .025. you will randomly select electric cables until you find a nondefective cable. what is the probability that you find that non defective cable on the sixth observation?

OpenStudy (kropot72):

Do you know how to find the z-scores for 8.025 and 7.975, these being the limits for a non-defective cable?

OpenStudy (anonymous):

kinda need help on that too

OpenStudy (kropot72):

\[z _{1}=\frac{X-\mu}{\sigma}=\frac{8.025-8.000}{0.02}=you\ can\ calculate\]

OpenStudy (anonymous):

now i get it... thank you so much.. took me like 2 hours to figure that out. :)

OpenStudy (kropot72):

Have you calculated the value of z1 yet?

OpenStudy (anonymous):

its 1.25

OpenStudy (kropot72):

Correct! Now for the z-score for 7.975: \[z _{2}=\frac{7.975-8.000}{0.02}=you\ can\ calculate\]

OpenStudy (anonymous):

-1.25

OpenStudy (kropot72):

Good work. Correct again! The next step is to find the probability that a randomly sampled cable has a diameter between 7.975 and 8.025. To do this we can use a standard normal distribution table and find the cumulative probabilities for z-scores of 1.25 and -1.25. Then the smaller probability is subtracted from the larger. You can use the standard normal distribution table here if you like: http://lilt.ilstu.edu/dasacke/eco148/ztable.htm

OpenStudy (anonymous):

2.50?

OpenStudy (kropot72):

Not really. You don't subtract the z-scores, rather you subtract the probability values which apply to the z-scores. These probability value are found from the table that I posted a link for. Have you looked at the table and found the probability for a z-score of -1.25?

OpenStudy (anonymous):

.1056?

OpenStudy (kropot72):

Very good work! Now can you find the probability for a z-score of 1.25?

OpenStudy (anonymous):

.8644

OpenStudy (kropot72):

The probability for a z-score of 1.25 is 0.8944 according to my reading of the table.

OpenStudy (anonymous):

oops you are correct. it was my mistake lol

OpenStudy (anonymous):

one question how come its 8.0 instead of 0.8?

OpenStudy (kropot72):

Therefore the probability that a randomly sampled cable has a diameter between 7.975 and 8.025 is given by: \[P(z _{1})-P(z _{2})=0.8944-0.1056=0.7888\] The probability that a randomly sampled cable does not have a diameter between 7.975 and 8.025 is given by: 1.0000 - 0.7888 = 0.2112 To find the required probability we need to find the probability that the 5 first samples result in a non-conforming cable each time and the sixth sample is a conforming cable. This is given by: \[P(non-defective\ is\ sixth\ observatio)=(0.2112)^{5}\times0.7888=you\ can\ calculate\]

OpenStudy (anonymous):

0.00033?

OpenStudy (kropot72):

Good work again! That is the correct answer :)

OpenStudy (anonymous):

one question: why we use mean 8 instead of 0.8?

OpenStudy (kropot72):

Very sorry! My bad :( I incorrectly read the question and took the mean diameter to be 8 instead of the stated 0.8. I will recalculate. Please wait.

OpenStudy (anonymous):

okay thank you :)

OpenStudy (kropot72):

When the correct value of mean is used, the z-scores come to the same values as with a mean of 8. Therefore all the calculations following the z-scores are still correct:)

OpenStudy (anonymous):

oh cool. thank you for helping me :D

OpenStudy (kropot72):

\[z _{1}=\frac{0.825-0.8}{0.02}=1.25\] \[z _{2}=\frac{0.775-0.8}{0.02}=-1.25\]

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

lol very close call.. same answer. once again thank you :)

OpenStudy (kropot72):

np :)